OpenStudy (anonymous):

Please solve this step by step. You dont have to explain it in detail. (16^1/4)^3

OpenStudy (anonymous):

\[(16^{1/4})^{3}\] Sorry I dont know how to write fractions properly.

OpenStudy (anonymous):

\[16^{1/4}*16^{1/4}*16^{1/4}=16^{1/4*1/4*1/4}=16^{1/4*4*4}=16^{1/64}\]

OpenStudy (anonymous):

\[(\sqrt[4]{16})^{3}\]